# Computation of attractor dimension and maximal sums of Lyapunov exponents using polynomial optimization

### v1

@@ Appendix A Exterior products and compound matrices
$\left(QAQ^{-1}\right)^{[k]}$ $=\frac{\mathrm{d}}{\mathrm{d}\delta}\left.\left(I+\delta QAQ^{-1}\right)^{(k)}\right|_{\delta=0}$
$=\frac{\mathrm{d}}{\mathrm{d}\delta}\left.\left(Q\left(I+\delta A\right)Q^{-1}\right)^{(k)}\right|_{\delta=0}$
$=\frac{\mathrm{d}}{\mathrm{d}\delta}\left.Q^{(k)}\left(I+\delta A\right)^{(k)}\left(Q^{(k)}\right)^{-1}\right|_{\delta=0}$
where the last equality uses properties 1 and 2 from proposition A.2. ∎
Multiplicative compound matrices are straightforward to compute directly, at least when $n$ is small. Additive compound matrices can be computed with symbolic manipulation. In our computational examples we compute additive compounds using Julia with the DynamicPolynomials.jl package [21]. Explicit formulas for the elements of additive compounds also exist [27].
We now have the ingredients to prove 15:
Lemma A.4.
If
$\frac{\mathrm{d}}{\mathrm{d}t}y_{i}(t)=Df(x(t))\,y_{i}(t)$ (126)
for $i=1,\dots,k$, and we define
$y^{k}(t)=y_{1}(t)\wedge\dots\wedge y_{k}(t),$ (127)
then
$\frac{\mathrm{d}}{\mathrm{d}t}y^{k}(t)=Df^{[k]}(x(t))\,y^{k}(t).$ (128)
@@
We have
$y_{i}(t)=D\phi^{t}(x_{0})\,y_{i}(0),$ (129)
so, using proposition A.2,
$y^{k}(t)=\left(D\phi^{t}(x_{0})\,y_{1}(0)\right)\wedge\dots\wedge\left(D\phi^{t}(x_{0})\,y_{k}(0)\right)=\left(D\phi^{t}x_{0}\right)^{(k)}y^{k}(0).$ (130)
Observe that $\phi^{0}(x_{0})=x_{0}$, and $\phi^{t+\delta}(x_{0})=\phi^{\delta}(\phi^{t}(x_{0}))$, so that
$D\phi^{t+\delta}(x_{0})=D\phi^{\delta}(x(t))\,D\phi^{t}(x_{0}).$ (131)
Then by proposition A.2,
$\left[D\phi^{t+\delta}(x_{0})\right]^{(k)}=\left[D\phi^{\delta}(x(t))\right]^{(k)}\,\left[D\phi^{t}(x_{0})\right]^{(k)},$ (132)
so by the usual definition of the derivative,
$\frac{\mathrm{d}}{\mathrm{d}t}y^{k}$ $=\lim_{\delta\to 0}\left(\frac{\left(D\phi^{t+\delta}(x_{0})\right)^{(k)}y^{k}(0)-\left(D\phi^{t}(x_{0})\right)^{(k)}y^{k}(0)}{\delta}\right)$
$=\lim_{\delta\to 0}\left(\frac{\left[D\phi^{\delta}(x(t))\right]^{(k)}-I}{\delta}\right)y^{k}(t).$
Furthermore, we have
$\left[D\phi^{\delta}(x(t))\right]^{(k)}$ $=\left[I+\delta Df(x(t))+o(\delta)\right]^{(k)}$
and so
$\frac{\mathrm{d}}{\mathrm{d}t}y^{k}$ $=\lim_{\delta\to 0}\left(\frac{\left[I+\delta Df(x(t))+o(\delta)\right]^{(k)}-I}{\delta}\right)y^{k}(t)$
$=\frac{\mathrm{d}}{\mathrm{d}\delta}\left.\left[I+\delta Df(x(t))\right]^{(k)}\right|_{\delta=0}y^{k}(t)$
$=\left(Df(x(t))\right)^{[k]}y^{k}(t),$
by A.1. ∎
Appendix B A specialised Lyapunov theorem
Here we present a Lyapunov theorem for the stability of one variable $w\in\mathbb{R}^{m}$ whose linear dynamics is controlled by a second variable $x$. This is applicable to both proposition 2.11 and proposition 2.12. The second part, whose proof follows that of similar results in Khalil [17], is the so-called converse Lyapunov theorem, which states that under stronger conditions, it is always possible to find a Lyapunov function and that this $V$ is quadratic in $w$.
Lemma B.1.
For a domain $\mathcal{B}\subset\mathbb{R}^{n}$, let $x(t)$ and $w(t)$ solve the ODE system
on $\mathcal{B}\times\mathbb{R}^{m}$, with $f\in C^{0}(\mathcal{B},\mathbb{R}^{n})$ and $A\in C^{0}(\mathcal{B},\mathbb{R}^{m\times m})$.
  1. 1.
    If there exists $V\in C^{1}\left(\mathcal{B}\times\mathbb{R}^{m}\right)$ such that
    for all $(x,w)\in\mathcal{B}\times\mathbb{R}^{m}$, then $w(t)$ is bounded forward in time for each $w(0)\in\mathbb{R}^{m}$.
  2. 2.
    If $A\in C^{1}(\mathcal{B},\mathbb{R}^{m\times m})$ is bounded and $f\in C^{1}(\mathcal{B},\mathbb{R}^{n})$, and if there exist $K>0$ and $\lambda>0$ such that
    $|w(t)|\leq Ke^{-\lambda t}|w(0)|$ (135)
    for all solutions to 133, then there exists $U\in C^{1}\left(\mathcal{B},\mathbb{R}^{m\times m}\right)$ such that 134 is satisfied with $V(x,w)=w^{\mathsf{T}}U(x)w$.
Proof.
For the first part, let $V\in C^{1}$ satisfy the two inequality conditions. For contradiction, suppose that there is a solution with $|w(t)|$ unbounded. The first $V$ condition then implies that $V(x(t),w(t))$ is unbounded forward in time. The second $V$ condition implies
@@ Appendix B A specialised Lyapunov theorem
$V(x_{0},w_{0})=\frac{2L}{1-e^{-2LT}}\int_{0}^{T}|w(t)|^{2}\,dt,$ (137)
where $(x(t),w(t))$ evolves under (133) from initial condition $(x(0),w(0))=(x_{0},w_{0})$, and $L$ is an upper bound for the operator norm $\|A(x)\|$ over $x\in\mathcal{B}$. Observe that, by linearity of $w(t)$ with respect to $w_{0}$, this is quadratic in $w_{0}$ and thus can be written
@@
$\leq 2L|w|^{2}$
i.e. $\frac{\mathrm{d}}{\mathrm{d}t}|w|^{2}\geq-2L|w|^{2}$. Thus Grönwall’s inequality gives us
$|w(t)|^{2}\geq e^{-2Lt}|w_{0}|^{2}$ (140)
@@
$V(x_{0},w_{0})\geq\frac{2L}{1-e^{-2LT}}\int_{0}^{T}e^{-2Lt}|w_{0}|^{2}dt=|w_{0}|^{2},$ (141)
satisfying the first condition of 134.
For the second condition of 134, we define
@@
$\begin{split}f(x(t))&\cdot D_{x}V(x(t),w(t))+\left[A(x(t))w(t)\right]\cdot D_{w}V(x(t),w(t))\\ &=\frac{\mathrm{d}v}{\mathrm{d}t}\\ &=\frac{2L}{1-e^{-2LT}}\left(|w(t+T)|^{2}-|w(t)|^{2}\right)\\ &\leq\frac{2L}{1-e^{-2LT}}\left(K^{2}e^{-2\lambda T}-1\right)|w(t)|^{2},\end{split}$ (143)
so if $T\geq\frac{1}{\lambda}\log{K}$ the condition is satisfied. ∎

### v2

@@ Appendix A Exterior products and compound matrices
$\left(BAB^{-1}\right)^{[k]}$ $=\frac{\mathrm{d}}{\mathrm{d}\delta}\left.\left(I+\delta BAB^{-1}\right)^{(k)}\right|_{\delta=0}$
$=\frac{\mathrm{d}}{\mathrm{d}\delta}\left.\left[B\left(I+\delta A\right)B^{-1}\right]^{(k)}\right|_{\delta=0}$
$=B^{(k)}\frac{\mathrm{d}}{\mathrm{d}\delta}\left.\left(I+\delta A\right)^{(k)}\right|_{\delta=0}\big(B^{-1}\big)^{(k)}$
$=B^{(k)}A^{[k]}\big(B^{(k)}\big)^{-1},$
where the third equality uses property 1 from proposition A.2, and the fourth equality uses property 3. ∎
Multiplicative compound matrices are straightforward to compute directly, at least when $n$ is small. Additive compound matrices can be found with symbolic computation.
Appendix B A specialized Lyapunov theorem
Here we present a Lyapunov function lemma on stability of $w(t)\in\mathbb{R}^{m}$ whose ODE is linear in $w$ but depends generally on $x(t)$ whose evolution can be nonlinear. This lemma is used to prove propositions 2.11 and 2.13. The second part, whose proof follows similar proofs in Khalil [18], is a so-called converse Lyapunov theorem that guarantees existence of a Lyapunov function under a sufficiently strong stability assumption.
Lemma B.1.
Let $x(t)$ and $w(t)$ solve the ODE system
on $\mathcal{B}\times\mathbb{R}^{m}$ with initial condition $(x_{0},w_{0})$, where $f\in C^{0}(\mathcal{B},\mathbb{R}^{n})$ and $A\in C^{0}(\mathcal{B},\mathbb{R}^{m\times m})$. Assume $\mathcal{B}$ is forward invariant for the $x$ dynamics.
  1. 1.
    Supopose each $x(t)$ is bounded for $t\geq 0$. If there exists $V\in C^{1}\left(\mathcal{B}\times\mathbb{R}^{m}\right)$ such that
    for all $(x,w)\in\mathcal{B}\times\mathbb{R}^{m}$, then each $w(t)$ is bounded for $t\geq 0$.
  2. 2.
    Suppose $\mathcal{B}$ is compact, $f\in C^{1}$ and $A\in C^{1}$. If there exist $K>0$ and $\lambda>0$ (uniform in $x_{0}$) such that
    $|w(t)|\leq Ke^{-\lambda t}|w_{0}|$ (126)
    for all solutions to 124, then there exists $U\in C^{1}\left(\mathcal{B},\mathbb{R}^{m\times m}\right)$ such that 125 is satisfied with $V(x,w)=w^{\mathsf{T}}U(x)w$.
@@ Appendix B A specialized Lyapunov theorem
For the first part, let $V\in C^{1}$ satisfy 125. For contradiction, suppose that there is a solution to 124 with $|w(t)|$ unbounded. The first condition on $V$ then implies that $V(x(t),w(t))$ is unbounded forward in time. The second condition on $V$ implies that
@@
$V(x_{0},w_{0})=\frac{2L}{1-e^{-2LT}}\int_{0}^{T}|w(t)|^{2}\,{\rm d}t,$ (128)
where $(x(t),w(t))$ evolves under (124) from initial condition $(x_{0},w_{0})$, and where $L$ is any value large enough that
$\max_{x\in\mathcal{B}}\|A(x)\|\leq L,$ (129)
where $\|\cdot\|$ denotes the $\ell^{2}$ operator norm, which is equal to the largest singular value of $A(x)$. This condition on $L$ is needed below to ensure $V(x_{0},w_{0})\geq|w_{0}|^{2}$. By linearity of $w(t)$ with respect to $w_{0}$, the expression 128 for $V$ is quadratic in $w_{0}$ and thus can be written
@@
$\leq 2L|w|^{2},$
so $\frac{\mathrm{d}}{\mathrm{d}t}|w|^{2}\geq-2L|w|^{2}$. Grönwall’s inequality then gives
$|w(t)|^{2}\geq e^{-2Lt}|w_{0}|^{2},$ (132)
@@
$V(x_{0},w_{0})\geq\frac{2L}{1-e^{-2LT}}\int_{0}^{T}e^{-2Lt}|w_{0}|^{2}{\rm d}t=|w_{0}|^{2},$ (133)
meaning that $V$ satisfies the first condition of 125.
For the second condition of 125, we define
@@
$\begin{split}f(x(t))&\cdot D_{x}V(x(t),w(t))+\left[A(x(t))w(t)\right]\cdot D_{w}V(x(t),w(t))\\ &=\frac{\mathrm{d}v}{\mathrm{d}t}\\ &=\frac{2L}{1-e^{-2LT}}\left(|w(t+T)|^{2}-|w(t)|^{2}\right)\\ &\leq\frac{2L}{1-e^{-2LT}}\left(K^{2}e^{-2\lambda T}-1\right)|w(t)|^{2}.\end{split}$ (135)
Choosing any $T\geq\frac{1}{\lambda}\log{K}$ ensures that the last line above is nonpositive, and so the second condition of 125 is satisfied. ∎